Operations Report – April 20th

Crew 315 Operations Report 20-04-2025

SOL: 0

Name of person filing report: Michael Andrews

Non-nominal systems: Bathroom sink drain was noticed to slowly drain

Notes on non-nominal systems: Plan to diagnose sink drain first thing on Sol 1

ROVERS

Spirit rover used:

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used:

Hours: 215.6 (1.3 hours used)

Beginning charge: 100%

Ending charge: 99%

Currently charging: No

Curiosity rover used:

Hours: 308.9 (1.3 hours used)

Beginning charge: 100%

Ending charge: 99%

Currently charging: No

Perseverance rover used:

Hours: 322.2 (1.3 hours used)

Beginning charge: 100%

Ending charge: 99%

Currently charging: Yes

General notes on rovers: Familiarized the crew with rover operation prior to simulation start.

Summary of Hab operations: This was our first day on Mars! Crew 315 arrived at MDRS around 1300, unpacked their belongings, and received their basic station training.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 2040
1) Per formula (see notes): 42.48 gallons (incorrect as this is the first measurement
2) Smart Home Dashboard (see notes): 13.12 gallons

Water (static tank, remaining gallons): 507.52 gallons (per formula)

Static tank pipe heater (on or off): No

Static tank heater (on or off): No

Toilet tank emptied (no or yes): No

Summary of internet: Completing reports, learning languages, communicating with family, completing professional tasks from work before simulation.

Summary of suits and radios: None used today

Summary of GreenHab operations: Familiarization with crew and GreenHab operations, only watered plants in the afternoon

WATER USE: 9 gallons

Heater (On or Off): Yes

Supplemental light (hours of operation): 1900 to 2100

Harvest (name, weight in grams): None

Summary of ScienceDome operations: No activity performed in dome today

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: Tape measurer used to measure water consumption

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: None

Operations Report – April 18th

Crew 314 Operations Report 18-04-2025

SOL: 12

Name of person filing report: Antoine Dubois (Engineer)

Non-nominal systems :

Station tidy-up.

Last tunnel check-up.

Notes on non-nominal systems :

Station tidy-up : The various modules have been tidied up and cleaned. We started with the outer modules, then the upper deck of the HAB and finished with the lower deck.

A final check of the tunnel was carried out before our departure and everything was in order.

ROVERS

Spirit rover used: /

Opportunity rover used: /

Curiosity rover used: /

Perseverance rover used: /

General notes on rovers: /

Summary of Hab operations: We spent time cooking, socializing, closing our experiments, and finishing our individual reports. We also engaged in physical exercise to maintain our fitness. To conclude the day and the simulation, we held our regular team meeting to review our progress and plan our departure tomorrow.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 8 pm
1) Per formula (see notes): 21.56 gallons
2) Smart Home Dashboard (see notes): 26.98 gallons

Water (static tank, remaining gallons): 229.79 gallons (per formula) & 233.44 gallons (Smart Home Dashboard)

Static tank pipe heater (on or off): no

Static tank heater (on or off) : no

Toilet tank emptied (no or yes): yes

Summary of internet: Research for experiments, mail, languages learning and music listening.

Summary of suits and radios: /

Summary of GreenHab operations: put notes here

I watered all the plants with water.

I cleaned the floor of the GreenHab and the shelves.

WATER USE: 8 gallons

Heater (On or Off): Heater on (setup temperature = 72 °F)

Supplemental light (hours of operation): 7-10 pm

Harvest (name, weight in grams): 9.3 g of mint, 14 g of chives, 30 g of spring onions, 16 g of parsley, 5 g of basil, 43 g of rainbow cherries

Summary of ScienceDome operations: put notes here

Observation of the viability and germination of AMF spores

Observation of the association between plantain roots and AMF spores

Cleaning of the ScienceDome

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: No work has been done, but we have replaced some tools that were not previously stored (like a wire cutters and a tape measure).

Summary of any observatory issues: No observation today due to bad weather.

Summary of health and safety issues: Everyone was in good health today, with no issues reported. There were no safety concerns to note.

Questions, concerns and requests to Mission Support: None

Crew 315 HSO Beginning Of Mission Checklist 21Apr2025

[title HSO Beginning Of Mission Checklist- April 21st]

Submitted by: Urban Koi, HSO

Crew: 315

Date: 20 April 2025

Part 1: Confirm Emergency Escape Routes Clear + Functional

  1. Stairs (Between Lower End Upper Deck): Clear + Functional

  2. Emergency window (Upper Deck, East Side): Clear + Functional

  3. Commander’s Window (Located in Commander’s Crew Quarter): Clear + Functional

Part 2: Inventory First Aid Kit

Main First Aid Cabinet (Medical Supplies + Medications)

  • Blood Pressure Monitor (Batteries OK)

  • Cotton Swabs (1.5 Boxes)

  • Medical Alcohol 91% (⅓ One Bottle, ⅔ One Bottle)

  • Antiseptic – Hydrogen Peroxide (Full One Bottle, ½ One Bottle)

  • Dental Floss (1 Nearly Full Unit)

  • Band-Aids (2 Boxes)

  • Sanitary Napkins (5 Pads)

  • Ibuprofen 200 mg (20/50)

  • Cutaneous Thermometer (Batteries OK)

  • FFP2 Masks (12 Masks)

  • Standard Surgical Masks (12 Masks)

  • Emergency Blankets (4 Units)

  • Emergency Splint (1 Unit)

  • Medical Tape (3 Incomplete Rolls)

  • Various types of Bandages (7 Rolls)

  • Motion Sickness Medication (1 Unit)

  • Nitrile Gloves (1 Pair)

  • Instant Cold Compress (Unused)

Small First Aid Kit (Medical Supplies + Medications)

  • Gauze Pads (10 Units)

  • Scissors (1 Unit)

  • Tweezers (1 Unit)

  • Antiseptic Towelettes (10 Units)

  • Triangular Bandage (1 Unit)

  • CPR Face Shield (1 Unit)

  • Medical Tape (1 Roll)

  • Nitrile Gloves (2 Pairs)

  • Burn Cream (3/10)

  • Hand Sanitizer (5/10)

  • Antibiotic Ointments (5/10)

  • Plastic Bandages (15/25)

  • Sterile Gel for Burns (1 Unit)

  • Eye Care: (2 Eye Patches, 1 Eye Wash)

  • Instant Cold Compress (1 Unit)

Part 3. Health + Environment Safety

Note Any Safety Issues: None

Note Any Health/Environmental Issues: None

Note Any Missing or Recommended Health and Safety Supplies: Recommend AED + EpiPen

Part 4. See Attached (Screen Capture of Equipment Table)

Operations Report – April 21st

Crew 315 Operations Report 21-04-2025

SOL: 1

Name of person filing report: Michael Andrews

Non-nominal systems: Furnace, Suit #7 Charger, Greenhouse Scale

Notes on non-nominal systems:

1. Oil leak on the main generator, which requires us to use the backup generator at night. This is being worked solely by Mission Support and does not require crew engineering.

2. It was noted that Suit #7 wasn’t charging, but further diagnosis from the Commander has isolated the issue to the #7 charger. Troubleshooting this will take place over the next day.

3. The greenhouse scale does not turn on, suspected batteries are dead and need to be replaced.

ROVERS

Spirit rover used: No

Hours: Not Used

Beginning charge: Not Used

Ending charge: Not Used

Currently charging: No

Opportunity rover used: Yes

Hours: 217.9 (0.2 hours used)

Beginning charge: 100%

Ending charge: 95%

Currently charging: No

Curiosity rover used: Yes

Hours: 311.5 (0.3 hours used)

Beginning charge: 99%

Ending charge: 94%

Currently charging: No

Perseverance rover used: Yes

Hours: 322.4 (0.2 hours used)

Beginning charge: 100%

Ending charge: 98%

Currently charging: No

General notes on rovers: None

Summary of Hab operations: The crew spent the morning trying out their project hardware (drones and Starlink mini) before starting simulation at noon. The crew split into two EVA teams and went to Marble Ridge as a MDRS tradition. Afterwards, we performed a 100cameras module with HSO Koi before dinner.

Water Use

Time of measurements: 2040
1) Per formula: 45.75 gallons
2) Smart Home Dashboard: Unable to calculate with main generator down

Water (static tank, remaining gallons): 461.78 gallons

Static tank pipe heater (on or off): No

Static tank heater (on or off): No

Toilet tank emptied (no or yes): No

Summary of internet: Used to complete reports, learn languages, share team photos via Google drive, streaming shows

Summary of suits and radios: Two EVAs were performed today, splitting the teams into two halves in order to complete the marble ritual.

· EVA 1: Suits 6, 8, 10 and Radios 1, 3, 4

· EVA 2: Suits 7, 9 and Radios 1, 2

Summary of GreenHab operations: Watered all plants and harvested some basil to go with olive oil for dipping with bread.

WATER USE: 10.8 gallons

Heater (On or Off): On

Supplemental light (hours of operation): 1900-2200

Harvest (name, weight in grams): Basil, Parsely, and Chives… all amounted to a handful but unable to weigh due to inoperable scale.

Summary of ScienceDome operations: Set up engineering project hardware (EVA Connectivity Kit and 3D Scanning Kit) and tested Starlink Mini outside at the observatory.

· EVA Connectivity Kit: Starlink took approximately 5 minutes to connect, and immediately uninstalled a software update. Crew Artist, Crew Journalist, Crew Engineer, and HSO were all able to connect successfully. Several speed tests were performed, with a maximum download of 186 Mbps and upload of 19.0 Mbps. 16% of the battery pack was consumed over 20 minutes.

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: (Tools used, work done)

· Leaking Sink: needle nose pliers, adjustable wrench, and plumber’s tape taken – all returned to the RAM. A little troubleshooting revealed that the stopper was clogged and needed to be removed, cleaned, and reinstalled. Plumber’s tape was added to the stopper’s side port to prevent leaking. Previously the stopper was wrapped externally with tape, but that was removed after the plumber’s tape solved the problem.

· Loose Staircase Mat: adjustable wrench, 3/8” deep socket, 2 nuts, and 2 washers were taken – nuts and washers were consumed, all other tools returned. The topmost staircase’s mats are loose, and it was discovered that no washer and nut was installed on the underside of the mat in two locations. A washer and nut was added, tightened, and now the stair provides less of a tripping hazard.

· Zip Ties: Only several open grommets were noted, mainly ones underneath other tarps. However, afternoon gusts liberated several zip ties that need replacing. I will make a routine of replacing the zip ties every morning before afternoon gusts arrive. In short, no zip ties were used today.

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: None

Operations Report – April 22nd

Crew 315 Operations Report 22-04-2025

SOL: 2

Name of person filing report: Michael Andrews

Non-nominal systems: Main Generator, Spacesuit Charger #7, Perseverance Rover

Notes on non-nominal systems:

1. (Unresolved from yesterday) Oil leak on the main generator, which requires us to use the backup generator at night. This is being worked solely by Mission Support and does not require crew engineering.

2. (Unresolved from yesterday) It was noted that Suit #7 wasn’t charging, but further diagnosis from the Commander has isolated the issue to the #7 charger. Troubleshooting this will take place over the next day.

3. The Perseverance Rover has two minor issues: the parking brake does not fully engage and have much effect, and the ignition block rotates fully if you turn the key too hard by accident.

ROVERS

Spirit rover used: No

Hours: Not Used

Beginning charge: Not Used

Ending charge: Not Used

Currently charging: Not Used

Opportunity rover used:

Hours: 219.0 hours (1.1 hours used)

Beginning charge: 100%

Ending charge: 64%

Currently charging: No

Curiosity rover used: Yes

Hours: 311.8 hours (0.5 hours used)

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Perseverance rover used: Yes

Hours: 323.8 hours (1.4 hours used)

Beginning charge: 100%

Ending charge: 89%

Currently charging: No

General notes on rovers: None

Summary of Hab operations: Two EVAs were performed today, the first to collect rock samples and the second to take desiccated soil samples. A 100cameras module on camera features was performed with the crew after breakfast. A major repair of the tunnels’ zip ties was also performed (details under RAM notes).

Water Use

Time of measurements (collect data for the same time window): 1930
1) Per formula (see notes): 29.41
2) Smart Home Dashboard (see notes): 25.25

Water (static tank, remaining gallons): 432.37

Static tank pipe heater: Off

Static tank heater: Off

Toilet tank emptied: No

Summary of internet: Used to complete reports, learn languages, share team photos via Google drive, researching projects

Summary of suits and radios:

· EVA 1: Suits 2, 3 and Radios 2, 3

· EVA 2: Suits 7, 8, 9 and Radios 1, 3, 5

Summary of GreenHab operations: Plants were watered, some basil leaves were harvested for meals. Since the fan doesn’t reliably work due to the main generator being inoperable, the temperature increased quicker than normal and needed to be regulated via opening the GreenHab door.

WATER USE: 10.4 gallons

Heater (On or Off): On

Supplemental light: 1900 – 2200

Harvest (name, weight in grams): Handful of basil leaves (unable to weigh due to faulty scale)

Summary of ScienceDome operations: Began the 3D scanning project with the rock samples recovered on EVA #3. A Rubik’s cube was first used as a simple model to scan, which didn’t work initially. I performed some more research using some downloaded manuals and videos to better understand the software before starting with a first rock sample. The results were much better, yielding a 3D model of a sample after about 2 hours of work. The model was lacking color, but more work will be done to add color to the texture and reduce scanning time. A power study for the EVA Connectivity kit was also performed.

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: (Tools used, work done) put notes here

· Zip Ties: Snips and 44 zip ties were taken from the hab, all to repair the tunnels’ tarps. I’m uncertain if the large amount of zip ties that needed to be replaced was because of the last several days of no engineering support between crew 314 and 315, or if certain zip ties were intentionally not replaced. Approximately half of those zip ties were applied on grommets that didn’t have any nearby zip ties on the ground.

· GreenHab Scale: I hoped that the scale needed a simple AAA battery swap, but it proved to be more complicated. The screen does not turn on even with fresh batteries from the exercise bike. I need to investigate the interior components for any obvious issues.

· Spacesuit #1 Repair: After personally attempting to wear two-piece suit #1, Commander Laude discovered that the neck ring had been installed upside down. I used a screwdriver and pliers to uninstall, flip, and reinstall the neck ring for future use.

· Other hardware used: measuring tape for water usage measurements, and an extension cord and 5 specimen jars were taken from the RAM for ScienceDome projects to be returned at end of mission.

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support:

· Is there a way to download the data from the Home Assistant program? That will allow the GreenHab specialist and myself to more accurately record the average temperature and water usage, respectively. Attempting to download the data on the graph view takes us to a new Safari page that cannot be viewed.

· Is there any issue with using a scale in the ScienceDome to weigh harvested plants in the GreenHab? Or must they remain in their respective areas?

Operations Report – April 23rd

Crew 315 Operations Report 23-04-2025

SOL: 3

Name of person filing report: Michael Andrews

Non-nominal systems: Charger #7, Space Suit #1, #3, Perseverance Rover, Curiosity Rover

Notes on non-nominal systems:

1. (Unresolved from yesterday) It was noted that Suit #7 wasn’t charging, but further diagnosis from the Commander has isolated the issue to the #7 charger. Troubleshooting this will take place over the next day.

2. Space Suit #3’s neck ring cord broke and needs to be reinstalled. Its cord was cannibalized from Suit #1 to support an immediate EVA and will be fixed this evening. Suit #3’s oxygen cord was also cross-threaded upon install and is stuck. Some quick work with some elbow grease and adjustable pliers will fix it this evening.

3. (Unresolved from yesterday) The Perseverance Rover has two minor issues: the parking brake does not fully engage and have much effect, and the ignition block rotates fully if you turn the key too hard by accident.

4. The Curiosity Rover has a lagging battery indicator. On EVAs, its battery life percentage will drop by only a few points and register much later on the following EVA. This is believed to be fixed by watering the battery with DI water, an action for Mission Support.

ROVERS

Spirit rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: Yes

Hours: 325.1 hours (1.3 hours used)

Beginning charge: 100%

Ending charge: 71%

Currently charging: No

Curiosity rover used: Yes

Hours: 313.5 hours (1.7 hours used)

Beginning charge: 100%

Ending charge: 67%

Currently charging: No

Perseverance rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: No

General notes on rovers: I performed a power study on all three rovers based on 6 EVAs of data. Perseverance has the best battery life, and Curiosity performs equally with Opportunity. However, until Curiosity’s battery is watered, its battery level does not update real-time.

Summary of Hab operations: The crew performed another 100cameras module after breakfast, cooked bread for lunch and burritos for dinner, and performed two EVAs.

Water Use

Time of measurements: 1930
1) Per formula: 41.66 gallons
2) Smart Home Dashboard: 46.5 gallons

Water (static tank, remaining gallons): 390.71 gallons

Static tank pipe heater: Off

Static tank heater: Off

Toilet tank emptied: Yes

Summary of internet: Used to complete reports, learn languages, share team photos via Google drive, researching projects

Summary of suits and radios:

· EVA 5: Suits 8,9 and Radios 3, 4

· EVA 6: Suits 2, 7, 10 and Radios 1, 2, 5

Summary of GreenHab operations: Watered all crops, cultivated fungal project, and opened door when temperatures exceeded 94 F.

WATER USE: 11.0 gallons

Heater: Yes

Supplemental light: 0000 to 0900 and then 1900 to 2200 (light was on all night due to script needing to be rebooted)

Harvest (name, weight in grams): tomatoes (66g), chives (8g), parsley (5g)

Summary of ScienceDome operations: It was a late night last night to map the second rock sample with the 3D scanner, but it proved to be successful and very informative. This morning a third sample was scanned in about an hour… half the time of previous samples! The Crew Journalist also began her desiccated soil experiment in the ScienceDome this afternoon. The fan that cools the power converter was turned on to level 3.

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations:

· Space Suit Repairs: As mentioned above, neck ring cords were swapped on suits 3 and 1 to facilitate EVA 6.

· Crew Artist’s Crazy Straw: Gagnon requested the ability to drink from his water bottle on EVAs without taking off his helmet. I fashioned a clean rubber hose from the science lab into a drinking apparatus that tapes to the suit helmet. With a little bit of adjustment, it worked well on EVA 6. Once the mission is over, it will be cleaned and returned to the dome.

· Tunnel Zip Ties: Only one zip tie was missing after yesterday’s evening gusts! A more thorough check will be performed in the AM. One zip tie was taken for install in next daylight.

· GreenHab Scale: a scale was repurposed from the ScienceDome to measure GreenHab harvests. Still troubleshooting the original scale to see if it can be fixed in the RAM.

· Other Tools Uses: Additional jars (2) were taken for rock specimens. A tool bag, carabiner, and rope tape measurer were taken for the Crew Journalist’s project, and a Rock Hammer was taken for rock specimens. All will be kept in the hab for EVA use until mission duration.

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: Great work on fixing the main generator Mission Support! This will lower noise levels and allow us to monitor station health via Home Assistant more consistently.

Operations Report – April 24th

Crew 315 Operations Report 24-04-2025

SOL: 4

Name of person filing report: Michael Andrews

Non-nominal systems: Suit Charger #7, Catalytic Heater

Notes on non-nominal systems:

· (Unresolved from yesterday) It was noted that Suit #7 wasn’t charging, but further diagnosis from the Commander has isolated the issue to the #7 charger. I was able to observe that all chargers appear to work, so more formal diagnoses are required.

· The catalytic heater near the exercise bike in the habitat does not work. It’s an engineering project to see if this can get properly started during this mission.

ROVERS

Spirit rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: Yes

Hours: 221.2 hours (1.0 hours used)

Beginning charge: 100%

Ending charge: 34%

Currently charging: Yes

Curiosity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Perseverance rover used: Yes

Hours: 326.9 hours (1.4 hours used)

Beginning charge: 100%

Ending charge: 68%

Currently charging: Yes

General notes on rovers: Curiosity and Perseverance had their batteries watered, so the team is intentionally waiting 8 hours (overnight) before using them again. Perseverance’s ignition and parking brake have been repaired. Curiosity’s power indicator was checked but does not have an easy fix, so we may need to manually estimate its battery life.

Summary of Hab operations: The team cooked bread, pasta for dinner, performed two EVAs, and congratulated Koi on winning her NASA grant!

Water Use

Time of measurements: 1845
1) Per formula: 40.03 gallons
2) Smart Home Dashboard: 34.5 gallons

Water (static tank, remaining gallons): 350.68 gallons

Static tank pipe heater (on or off): Off

Static tank heater (on or off): Off

Toilet tank emptied (no or yes): No

Summary of internet: Used to complete reports, learn languages, share team photos via Google drive, researching projects

Summary of suits and radios:

· EVA 7: Suits 7,9 and Radios 3, 4

· EVA 8: Suits 8, 10 and Radios 3, 4

Summary of GreenHab operations: Watered all plants, harvested herbs for dinner and bread.

WATER USE: 10.3 gallons

Heater (On or Off): On

Supplemental light (hours of operation): 1900-2200

Harvest (name, weight in grams): harvested 20 combined grams of chives, basil, and parsley.

Summary of ScienceDome operations: None

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations:

· Tunnel Zip Tie Repair: Only one zip tie was needed to be replaced on the tunnels after

· Suit Repair: Suit #1’s repair of the neck cable ring was extremely simple once Scott Davis from Mission Support informed me of the extra space suit pieces. I had tried to fabricate my own cable ring with cable from the RAM and crimping it with no success.

· GreenHab Scale: For reasons unknown, fully disassembling the Greenhab scale and reassembling it worked! I guess that there was some dirt between the circuit board and the buttons that could have prevented contact. The ScienceDome scale was returned to the ScienceDome after the GreenHab scale was returned to the GreenHab.

· 3D Scanning Shroud: I was able to craft a rudimentary shroud for en situ 3D scanning for my project, but I want to confirm that all components can be sent out on EVA on Sol 6 or 7. Please see questions to Mission Support below.

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: Two requests

· We would like to see if we can get the closeout checklist for the season early. We want to start reviewing the actions needed to close up MDRS for the summer.

· For my 3D scanning project, I plan to scan rock samples on EVA and need a way to create a large area of shade for the activity. I was able to create a shroud with the two sawhorses, clamps, and the fire blanket from the RAM. As soon as I tested the concept, I returned the blanket to its location. However, I do want to see if I can take these items with me on future EVAs? I would only take the RAM’s blanket, would ensure other crew members aren’t in the RAM when there is no blanket, and would call out the blanket’s use on EVA. If this is approved, let me know what is necessary to properly communicate its use.

Operations Report – April 25th

Crew 315 Operations Report 25-04-2025

SOL: 5

Name of person filing report: Michael Andrews

Non-nominal systems: Suit Charger #7, Catalytic Heater, ScienceDome Radio Charger

Notes on non-nominal systems:

· (Unresolved from yesterday) It was noted that Suit #7 wasn’t charging, but further diagnosis from the Commander has isolated the issue to the #7 charger. I was able to observe that all chargers appear to work, so more formal diagnoses are required.

· (Unresolved from yesterday) The catalytic heater near the exercise bike in the habitat does not work. It’s an engineering project to see if this can get properly started during this mission.

· The ScienceDome radio chargers do not work. There seems to be a loose component internally, because the radios will flicker on and off between charging. Furthermore, docking the ScienceDome radios in the GreenHab hold a charge nominally. Radios were swapped between the two structures, and I will disassemble the charger tomorrow.

ROVERS

Spirit rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Curiosity rover used: Yes

Hours: 313.8 hours (0.3 hours used)

Beginning charge: 100%

Ending charge: 93%

Currently charging: No

Perseverance rover used: Yes

Hours: 328.5 hours (1.5 hours used)

Beginning charge: 100%

Ending charge: 55%

Currently charging: Yes

General notes on rovers: None

Summary of Hab operations: We performed two EVAs, made corn muffins with cheese and herbs, and pizza for dinner. The team also had a discussion about water conservation so we can ensure we complete the mission with our allotted 500 gallons. I also collected trash and replaced trash liners. Winds were very high in the afternoon, gusting above 30 mph. We also completed the 500 piece puzzle yesterday and have progressed to the 1000 piece puzzle.

Water Use

Time of measurements: 1910
1) Per formula: 22.87 gallons
2) Smart Home Dashboard: 25.5 gallons

Water (static tank, remaining gallons): 327.31 gallons

Static tank pipe heater: Off

Static tank heater: Off

Toilet tank emptied: No

Summary of internet: Used to complete reports, learn languages, share team photos via Google drive, researching projects

Summary of suits and radios:

· EVA 9: Suits 7,8 and Radios 2,3

· EVA 10: Suits 1, 9 and Radios 1, 2

Summary of GreenHab operations: Watered plants, harvested vegetables and herbs for lunch and dinner, and watered fungi project.

WATER USE: 10 gallons

Heater (On or Off): On

Supplemental light (hours of operation): 1900-2200

Harvest (name, weight in grams): 16 tomatoes, 32g basil, 2g parsley, 22g green onion

Summary of ScienceDome operations: Used the ScienceDome for additional rock scanning, including a multi-scan project for a shell collected from Sea of Shells. Noticed that the radio charger in the ScienceDome is not charging radios and needs troubleshooting.

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: No significant work in the RAM today – only grabbed one zip tie for the tunnel repair and loaded 1 bag of burn trash and 1 bag of no-burn trash in the RAM airlock for Mission Support to pick up.

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: One of the space suit helmets for the two-piece suits has a cracked neck ring. I discovered this helmet on the top shelf so it appears to have cracked a while ago. Do we have an extra red neck ring for the helmet that I can swap out?

Operations Report – April 7th

Crew 314 Operations Report 04-07-2025

SOL : 1

Name of person filing report : Béatrice Hollander (Crew Commander)

Non-nominal systems :

Notes on non-nominal systems :

ROVERS:

Spirit rover used: EVA #1

Hours: 272.6 hours (with 0.1 used)

Beginning charge : 100%

Ending charge : 95%

Currently charging : No.

Opportunity rover used : EVA #1

Hours : 212.4 hours (with 0.2 used)

Beginning charge : 100%

Ending charge : 93%

Currently charging : No

Curiosity rover used: EVA #2

Hours: 306.1 hours (0.2 used)

Beginning charge: 100

Ending charge: 95

Currently charging: No

Perseverance rover used: EVA #2

Hours: 317.2 hours (0.2 used)

Beginning charge: 100

Ending charge: 95

Currently charging :No

General notes on rovers : Nothing special on EVA #1 and EVA #2

Summary of Hab operations :

WATER USE: 64.04 gallons

Water (static tank): 485.96 gallons

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: no

Summary of internet: Used for music playing, mails, research for experiments, learning languages.

Summary of suits and radios:

Radios 1,2 and 5 were used on EVA #1

Light Suits 5, 6 and 9 were used on EVA #1

Radios 3,4,6 and 8 were used on EVA #2

Light Suits 2,4,5, and 6 were used on EVA #2

Summary of GreenHab operations: I watered all the plants (except dead plants) with the Miracle-Gro solution (1 cup for 2 gallons).

WATER USE: 4.8 gallons

Heater: Off

Supplemental light: N/A

Harvest: 3 g of mint and 5 g basilic

Summary of ScienceDome operations:

Dual split:

During the morning: preparation of all the materials (Spectral lamp, RPM (gravity), temperature sensor) for the stress experiment on arbuscular mycorhizal fungi.

During the afternoon: isolation of spores of arbuscular mycorrhizal fungi.

During the evening: end of isolation of spores and start of the experiment (stress application)

Milieu, petri dish preparations

Summary of RAM operations: (Tools used, work done)

A total of 24 zip ties were used to secure the tarps to the metal structure between each module. The ends of the zip ties were cut using a wire cutter. Additionally, duct tape was used to patch a large hole in the tarp near the Science Dome.

The crew engineer, Antoine, used a drill to attach the dust collector brackets to the stakes during the EVA (tools used: drill, mete r& 6 zip ties).

Summary of any observatory issues: Nothing as no observations done yet

Summary of health and safety issues : Our engineer was still a bit blurry this morning, but he was still able to participate in the daily tasks. We will take an eye on his health issue for the days to come, and keep you updated. The other members were feeling fine.
First Aid Kits : All medical supplies have been filled.

Questions, concerns and requests to Mission Support :

Request to Mission Support :

EXPERIMENT : The team has applied for support to collect a parcel containing important samples for our journalist’s experiment.

HSO :

As mentioned in the email’s response (HSO pre-mission checklist) : We have two issues with the walkie-talkies from SciDome: the one working on channel 10 is dead and doesn’t work on any other charging base. Unfortunately, the charging base for the SciDome is also no longer working.

As mentioned in the email’s response (HSO pre-mission checklist) : there’s no new alcaline battery left in the drawer (precicely : Alcaline Battery 9V 522-6LF22 -6AM6). We have received a response from Mission Support that batteries will be coming in the beginning of next week.

Operations Report – April 8th

Crew 314 Operations Report 08Apr2025

SOL:2

Name of person filing report: Antoine Dubois (Crew Engineer)

Non-nominal systems :

Spacesuit #2 has been repaired

Spacesuit #3 is not yet fully repaired

We suspect a contact failure with headset #2

The CO detector in the science dome is not working

Water leak toilet

Notes on non-nominal systems :

Spacesuit #2: The support frame is now correctly attached to the spacesuit backpack.

Spacesuit #3: The intermediate octagonal piece connecting the backpack to the suit hose is missing. However, we do have the screw thread to connect it to the suit.

Headset #2: We’ll take a closer look tomorrow before confirming whether it’s damaged.

Water leak toilet : Connection between water inlet pipe and bowl has broken.

ROVERS

Spirit rover used: EVA #3

Hours: 272.7

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging) 95%

Currently charging: NO

Opportunity rover used:EVA #3

Hours: 212.6

Beginning charge:, 100%

Ending charge: 95%

Currently charging: NO

Curiosity rover used: EVA #4

Hours: 306.5

Beginning charge: 100%

Ending charge: 90%

Currently charging: NO

Perseverance rover used: EVA #4

Hours: 317.9

Beginning charge: 100%

Ending charge: 89%

Currently charging: NO

General notes on rovers: none

Summary of Hab operations: Today in the Habitat, we focused on cooking, socializing, completing our individual reports, and working on our respective experiments. Additionally, we carried out the medical simulation as planned by Odile’s experiment. Finally, we held our usual team meeting to discuss progress and upcoming tasks.

Water Use (please use both methods to estimate water usage)

Time of measurements (collect data for the same time window): 8:00 p.m.
1) Per formula (see notes): 67.83 gallons
2) Smart Home Dashboard (see notes): 48.98 gallons

Water (static tank, remaining gallons): 418.83 gallons (per formula) & 436.97 gallons (Smart Home Dashboard)

Static tank pipe heater (on or off): on

Static tank heater (on or off) : on

Toilet tank emptied (no or yes): no

Summary of internet: used for music playing, mails, research for experiments, learning languages.

Summary of suits and radios:

EVA #3 : 7, 9 and 10

EVA #4 : 8, 9 and 10

Summary of GreenHab operations: Use of 300 g of substrate (in the wheelbarrow) for an experiment. I watered all the plants with water. I removed all the dead plants, threw them in the garbage provided for that purpose, and put all the soil from the dead plants’ pots in the wheelbarrow. I swept the floor of the Greenhab.

WATER USE: 7.5 gallons

Heater (On or Off): Heater on (setup temperature = 72°F)

Supplemental light (hours of operation): 7-10 pm

Harvest (name, weight in grams): None

Summary of ScienceDome operations:

Milieu preparation and assembling UV lamp

Overnight culture inoculation

Preparation of pots for growing plantain associated with arbuscular mycorrhizal fungi.

Preparation of different solutions required for experiment (plant root staining).

Solving a problem with the spectral lamp.

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations : Use of 6 zip ties, 1 meter of blue tape, 50 cm of wire for an experiment, 1 meter of tape to measure the tank’s depth, and pliers to repair the toilet leak. No actual operations were conducted within the RAM itself.

Summary of any observatory issues: /

Summary of health and safety issues : Our crew engineer is steadily improving each day. As for other issues, we encountered minor incidents, such as cuts, for which we applied band-aids to Batoul and Arnaud. Additionally, a minor issue occurred in the Science Dome when the autoclave overheated, causing a power failure and triggering the gas alarms due to the release of a large amount of steam. We had to request assistance from Mission Support to resolve the situation.
About the inventory : 4 band aids have been used today, and 1 antiseptic towelettes.

Questions, concerns and requests to Mission Support : /

Copyright © The Mars Society. All rights reserved. | Main Site